A 0.500 cm diameter plastic sphere, used in a static electricity demonstration, has a uniformly distributed 40.0 μC charge on its surface. What is the potential near its surface?


1
Expert's answer
2021-01-07T11:00:36-0500

By tribute of symmetry, the potential at the surface of the sphere will be the same as that of a point charge placed at the center of the sphere, having the same charge. That is, the radius would be 2.5 mm, as it is half the diameter, and the potential can then be calculated as:


"V=k\times \frac{q}{r}=9\times10^9\times\frac{4\times10^{-11} }{0.0025}=\ 144\ (V);"


Result: "144 \ V"